Please show your work. Thank you
Q4. Use the semi-annually compounded yield curve in the following table to price the some fixed income securities: 0.50 1.00 1.5 6.49% (a) 1.5-year zero coupon bond (b) 2-year coupon bond paying 15% semiannually (d) 1.5-year coupon bond paying 9% annually (e) 2-year floating rate bond with zero spread and semiannual payments (f) 1.5-year floating rate bond with zero spread and annual payments. For this question, assume ri (-0.5, 0.5)-6%. (h) 1.5-year floating rate bond with 40 basis point spread with annual payments. For this question, assume ri (-0.5, 0.5-696.
Answer #1

Sol (a) The price of zero coupon bond at 1.5 year will be calculated by using Present Value formula where we are assuming bond issued at par value i.e. $100.00 thus our Future Value = $100.00 the yield for 1.5 years is mentioned as 6.84% Time is 1.5 years thus Present Value = 100*(1/1.0684)^1.5= $90.55 as price of bond.

(b) The price of two year coupon bond paying 15% semiannually can be calculated using Present Value Annuity formula where annuity will be Coupon payment of assumed par value $100.00 bond as $7.50 in half year. The time will be 2N thus 4 and rate of yield is mentioned in the table as 6.88%. Thus price of bond will be $51.16 using excel formula of =PV(0.0688,4,7.50,-100,0).

(c) The price of 1.5 year coupon bond with 9% annual payment should be assuming par value as $100 N=1.5years, PMT=$9.00 and Rate= 6.84*2=13.68% PV(0.1368,1.5,9,-100,0)= $70.99

(d)